The molecular weight of methanol (CH 3 OH) a poisonous compound sometimes found moonshine,is a _____ ( rounded to one decimal place)
Answer #1

The molecular formula is CH3OH = CH4O

Atomic weight of Carbon = 12.0 g/mol

Atomic weight of Hydrogen = 1.0 g/mol

Atomic weight of Oxygen = 16.0 g/mol

Molecular weight = (1*12.0) + (4*1.0) + (1*16.0) = 12.0 + 4.0 + 16.0 = 32.0 g/mol

Molecular weight of Methanol = 32.0 g/mol